When 1.50 mol of CO2 and 1.50 mol H2 are placed in a 0.750-L container at 395ºC, the following

equilibrium is achieved: CO2(g) + H2(g) ↔ CO(g) + H2O(g).

If Kc = 0.802, what are the concentrations of each substance in the equilibrium mixture?

USE 3 SIG FIGS for each answer?

CO2 & H2 =

CO & H2O =
1
Expert's answer
2020-12-22T04:31:16-0500

Solution:

The balanced chemical equation:

CO2(g) + H2(g) ↔ CO(g) + H2O(g)

The equilibrium expression for the reaction system:



The initial concentration (Co) of each gas is:

Co(CO2) = n(CO2) / Vcontainer = 1.50 mol / 0.750 L = 2.00 mol/L = 2.00 M

Co(H2) = n(H2) / Vcontainer = 1.50 mol / 0.750 L = 2.00 mol/L = 2.00 M

Co(CO) = Co(H2O) = 0 (at the beginning of the reaction).


The reaction table is:



The equilibrium concentration of each gas is:

[CO2] = [H2] = 2.00 - x

[CO] = [H2O] = x


Hence,



or



x = 0.8955 * (2.00 - x)

1.11664x = 2.00 - x

2.11664x = 2.00

x = 0.9449 = 0.945


Thus:

[CO] = [H2O] = x = 0.945 M

[CO2] = [H2] = 2.00 - 0.945 = 1.055 M = 1.06 M


Answer:

[CO] = [H2O] = 0.945 M;

[CO2] = [H2] = 1.06 M.